# Susan Athey    Emil Palikot

### v4

@@ Abstract
Women remain underrepresented in technology despite many support programs, partly due to limited evidence on which interventions work and how to allocate scarce program slots. We evaluate two approaches to facilitating technology-sector transitions for women in Poland: Challenges, a scalable and cheap-to-operate (\$15/person) online program developing job-relevant portfolios (designed and implemented during this research), and one-on-one mentoring. The programs address distinct labor market barriers: expanding professional networks and teaching their effective use (Mentoring) versus providing credible skill signals (Challenges). Randomizing oversubscribed admissions, we find both programs substantially increase technology employment twelve months post-completion, Mentoring by 15 percentage points (p.p.) and Challenges by 11 p.p. Exploiting treatment effect heterogeneity, we show that algorithmic targeting based on predicted treatment effects achieves outcomes 86% larger than random assignment at current capacity levels. While mentors successfully identify high-benefit candidates (outcomes 48% above random selection), algorithmic allocation using observable characteristics yields an additional 11% improvement. The results demonstrate that (i) relatively low-cost interventions can meaningfully address gender gaps in technology, and (ii) data-driven allocation rules can substantially enhance program effectiveness even when expert judgment performs well. JEL: J24, J08, J16, C93
@@ 1 Introduction
An important concern for labor policymakers is identifying and implementing effective programs to support workers transitioning into high-growth sectors. The technology sector exemplifies this challenge: despite rapid employment growth and attractive wages, women hold only 30% of technology jobs worldwide.111unitednations2021 estimates the share of women in the technology sector at approximately 30%. Many programs exist to support women in technology, including NGO initiatives (Women in Technology International, Girls Who Code), corporate programs (Women in Tech from Salesforce, The Women at Microsoft), and government efforts (UN’s Girls in ICT Day, European Commission’s Women in Digital Scoreboard). While women face multiple barriers entering technology, this paper examines two specific labor market frictions. [^2] The gender gap in technology has been attributed to discrimination in hiring and promotion (reuben2014gender; moss-racusin2012science), work-family conflict and child penalties (goldin2014grand; kleven2019children; bertrand2010dynamics), and underrepresentation in STEM education pipelines (bayer2016diversity; carrell2010sex). We focus on network and information barriers because short-term labor market interventions can directly address these constraints, unlike structural factors requiring longer-term policy responses. First, women have weaker professional networks than men, limiting access to job opportunities (greguletz2019women; forret2004networking). Second, gender stereotypes lead employers to discount women’s technical abilities, creating information frictions that credible skill signals could reduce (cheryan2013stereotypical; del2022more; murciano2022missing). Understanding which constraints bind, and for whom, is critical for designing effective policy. If network barriers dominate, interventions should expand professional connections; if information frictions dominate, scalable credentialing programs may suffice. Alternatively, both frictions may bind simultaneously for different populations, necessitating multiple program types and optimal assignment rules. We provide experimental evidence distinguishing between these possibilities.
@@
We evaluate two interventions implemented in Poland by Dare IT, a social impact organization, each targeting one of these frictions. Mentoring pairs participants with senior female technology professionals for four months of one-on-one sessions. While sessions are personalized, mentors consistently emphasize networking: facilitating introductions to hiring managers, expanding mentees’ professional connections, and teaching effective relationship leverage. Challenges is a scalable online program (\$15 per participant) we designed during this research to address skill signaling barriers. The program helps participants build portfolios demonstrating technical proficiency through six bi-weekly industry-designed assignments, culminating in complete projects, mobile app designs or React applications, that participants reference in job applications. [^3] Program development began with structured interviews with hiring managers at twelve Polish technology firms, who uniformly cited the absence of credible practical experience signals in female applicants’ profiles. We designed Challenges to be delivered entirely online, centered on peer interaction rather than scarce mentors, and based on realistic business cases to ensure employer credibility. The program has been replicated across multiple specializations (software testing, cloud computing, UI design) and delivered in Ukrainian for refugees. We exploit both programs’ substantial oversubscription to randomize access and measure employment outcomes via LinkedIn profiles over twelve months. [^4] Technology jobs include positions in technology firms (excluding finance, legal, HR, accounting roles) or technical positions elsewhere (software development, IT support, data analytics). UX and front-end programming roles are classified as technology jobs regardless of firm type.
@@
Both programs generate large employment effects, but with markedly different patterns of treatment effect heterogeneity that reveal which constraints bind for different populations. Mentoring increases technology employment by 15 percentage points (SE 5.7, from a 46% control baseline); Challenges by 11 percentage points (SE 4.8, from 27%). Evidence supports distinct mechanisms. Mentoring increases LinkedIn connections by 52.7 (SE 20.8), a 23% increase, while Challenges shows no detectable network expansion.
@@
The programs benefit different populations, with heterogeneity patterns aligning with each mechanism. Professional experience reveals the starkest contrast: Challenges generates 28 percentage points larger effects for participants with under five years of experience compared to those with longer careers, while Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced workers. This indicates that early-career women benefit from portfolio signals that compensate for limited work history, while experienced women benefit from network expansion that leverages their accumulated professional standing. Geography provides additional evidence: Mentoring generates 26 percentage point effects for women in small towns compared to only 3 percentage points for Warsaw residents (difference: 23 percentage points, SE 14), while Challenges shows no significant geographic heterogeneity (difference: -2 percentage points, SE 12). This pattern is consistent with network barriers binding tightly outside major urban centers where local technology ecosystems are sparse, while online portfolio signals function equally regardless of location. Mentor characteristics also predict treatment effects: participants matched to mentors with managerial experience or over five years in technology show larger treatment effects, consistent with senior mentors possessing richer networks and better access to hiring decision-makers. These differential patterns demonstrate that both network barriers and information frictions constrain technology employment, but for distinct populations defined by career stage or geography.
@@
The differential treatment effects create an allocation problem when program capacity is limited. Both programs remain substantially oversubscribed (5:1 applicant-to-spot ratios for Mentoring; 300 spots filled in four hours for Challenges), requiring selection among applicants. We evaluate counterfactual assignment policies using the methods of athey2021policy and zhou2023offline, estimating which applicants should receive which program to maximize aggregate outcomes. A targeted policy assigning 15% of applicants to each program based on predicted treatment effects achieves outcomes 86% larger than random assignment. This gain reflects two sources of value: selecting individuals with highest treatment effects and matching individuals to their optimal program type. The matching component generates persistent value: even when capacity expands to 50% per program (universal coverage), targeted allocation maintains outcomes 23% larger than random assignment. This demonstrates that complementarity between individual characteristics and program mechanisms creates allocation gains beyond simple capacity expansion, treatment effect heterogeneity enables welfare-improving assignment even when most applicants can be served.
@@
We further compare algorithmic targeting to expert human judgment. Mentoring relies on experienced female technology professionals to select mentees from pre-screened applicants based on pre-recorded short videos and the assessment of the application materials (cover letter, resume, and registration survey). This expert selection substantially outperforms random allocation, achieving outcomes 48% larger. Yet algorithmic allocation using only six observable characteristics (education, location, age, professional experience, degree level, domain preference) achieves outcomes 63% larger than random assignment, an additional 11% improvement over mentor judgment. The two approaches, by experts and by the algorithm, achieve high-degree of concordance (70%): the algorithm largely validates expert choices while identifying additional high-benefit candidates. Despite mentors’ rich soft information, systematic patterns in treatment response remain that simple demographic targeting can exploit. The result extends recent work showing algorithms can complement expert decisions in domains from bail (kleinberg2018human) to medical diagnosis (mullainathan2022diagnosing), demonstrating that even sophisticated expert judgment leaves room for data-driven improvement when treatment effects vary systematically with observable characteristics.
@@ 2 Literature Review
This paper contributes to the literature on treatment effect heterogeneity and optimal policy assignment. Building on recent advances in policy learning using machine learning methods (athey2017econometrics; kitagawa2018should; athey2021policy), we examine whether data-driven targeting can improve program allocation decisions in active labor market policy. Theoretical work demonstrates that when treatment effects vary across individuals and program capacity is limited, policies that assign treatment based on predicted treatment effects can substantially improve welfare relative to random allocation or allocation based on baseline characteristics (kitagawa2018should; bhattacharya2012inferring). Empirical applications of these methods have demonstrated value across diverse domains: customer retention (ascarza2018retention), marketing (hitsch2018heterogeneous), development programs (haushofer2022targeting), agriculture (kasy2021adaptive), and medical treatment (inoue2023machine).
@@
We further contribute to the growing literature examining whether algorithmic decision-making can complement or improve upon expert human judgment. Recent work demonstrates that machine learning algorithms can match or exceed expert performance in domains from bail decisions (kleinberg2018human) to medical diagnosis (mullainathan2022diagnosing), often by identifying systematic patterns that experts miss or cannot consistently apply. Similar to findings in online labor markets where revealing information about worker quality improves matching outcomes (pallais2014inefficient), we show that observable characteristics predict treatment effects systematically.
@@
We demonstrate that data-driven targeting can improve upon expert human judgment. Mentor-based selection achieves treatment effects 48% larger than random allocation. Yet algorithmic targeting using only six observable characteristics achieves an additional 11% improvement. The 70% concordance between methods indicates complementarity: algorithms largely validate expert choices while identifying additional high-benefit candidates, demonstrating that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity.
@@
This paper contributes to the literature on mentoring effectiveness in improving career outcomes. Several influential papers focus on mentoring in academic contexts: dennehy2017female; resnjanskij2021can; gardiner2007show; ginther2020can; blau2010can show that mentorship increases sense of belonging among female engineering students, youth’ learning outcomes, and tenure rates, publications, and federal grants for women in research careers. [^5] The impact of mentoring on career outcomes has also been studied using observational data. ragins1999mentor survey over 600 individuals and find that mentored respondents have better career outcomes than non-mentored respondents, with effects varying by the gender composition of the mentoring pair. kammeyer2008quantitative conduct a meta-analysis concluding that mentoring improves job and career satisfaction. Mentoring programs may operate partly through expanding professional networks, which play a critical role in labor market outcomes (granovetter1973strength; ioannides2004job; pallais2016referential). We provide randomized trial evidence on mentoring effectiveness in labor market transitions outside academic settings, demonstrating large effects on technology employment for women career-changers. Crucially, we provide experimental evidence that mentoring causally expands professional networks: Mentoring increases participants’ LinkedIn connections by 52.7 (SE 20.8), a 23% increase from baseline, while the non-network-based Challenges program shows no statistically significant effect. We further examine the evaluative selection mechanism inherent in mentor-based programs, where mentors exercise discretion in choosing mentees, and demonstrate that this selection generates substantial treatment effect heterogeneity with observable characteristics of both mentors and mentees predicting treatment effects systematically.
@@
We contribute to the extensive literature evaluating active labor market training programs. card2018works conduct a comprehensive meta-analysis of over 200 studies, finding that traditional training programs aimed at unemployed workers exhibit near-zero short-term employment effects, with modest positive impacts emerging only 2-3 years post-completion. [^6] See also heckman1999economics, card2010active, and crepon2016active for reviews of ALMP effectiveness. Programs incorporating on-the-job skill development tend to outperform classroom-based alternatives (sianesi2008differential; lechner2000microeconometric), though even intensive apprenticeships show mixed results (biewen2014effectiveness). fein2018bridging show that YearUp, a sectoral training program targeting disadvantaged youth, generates 40% earnings gains. However, recent evidence suggests that programs targeting skill signaling and information frictions generate substantially larger employment gains than traditional classroom training (abebe2021anonymity; bassi2022screening; carranza2020job; athey2024moocs). Building on the theoretical work of spence1978job, this literature demonstrates that credentials demonstrating capability can significantly boost earnings (tyler2000estimating; hadavand2018can) and generate market-level benefits when employers face uncertainty about worker quality (pallais2014inefficient). Recent evidence further suggests that training effectiveness may depend on complementary interventions: batista2022closing find that business training for female microentrepreneurs generates significant profit increases only when combined with mobile money access, highlighting that releasing multiple constraints simultaneously may be necessary for training programs to translate into economic gains.
@@
We provide experimental evidence on training program effectiveness in technology sector transitions, a context involving higher-skilled, career-changing women rather than the unemployed workers typically studied. We document substantial short-term effects (visible within twelve months) for both programs, aligning with the signaling literature rather than the null findings in card2018works’ meta-analysis. We directly compare two program modalities: personalized one-on-one mentoring versus scalable online project-based training. This head-to-head comparison reveals complementarity rather than substitutability: the programs serve different populations and generate heterogeneous treatment effects that enable targeted assignment. Our subjects are women seeking technology roles who can utilize skill signaling to counter gender-based statistical discrimination.
@@
Finally, this paper adds to the literature on the gender gap in the technology sector. The persistent gender wage gap reflects both occupational segregation and within-occupation differences (juhn2017specialization). Various reasons for low women’s participation in the technology sector have been studied, including women’s tendency to under-report programming skills (murciano2022missing) and the impact of occupational stereotypes (cheryan2013stereotypical; del2022more). Strategies proposed to counter these barriers include increasing role model visibility (correll2016succeed) and shifting perceptions of technology roles. We contribute to this literature by demonstrating how mentoring and credentialing programs can enhance women’s success rates in obtaining technology jobs.

### v5

@@ Abstract
We evaluate two interventions facilitating technology-sector transitions for women in Poland: Mentoring, focused on expanding professional networks, and Challenges, focused on building credible skill signals. Randomizing oversubscribed admissions, we find both programs substantially increase technology employment at twelve months—by 15 percentage points for Mentoring and 11 p.p. for Challenges. The distinct mechanisms through which the programs operate translate to heterogeneous treatment effects across geography, career stage, and baseline credentials. These differential effects create scope for improved allocation: algorithmic targeting across programs outperforms random assignment by 86% and experts’ selection into Mentoring by 11%. JEL: J24, J08, J16, C93
@@ 1 Introduction
Facilitating worker transitions into high-growth sectors is a core challenge for labor market policy, particularly for underrepresented groups. The technology sector is a prime example: despite rapid employment growth and attractive wages, women hold only 30% of technology jobs worldwide. [^1] United Nations (2021) estimates the share of women in the technology sector at approximately 30%. Many programs exist to support women in technology, including NGO initiatives (Women in Technology International, Girls Who Code), corporate programs (Women in Tech from Salesforce, The Women at Microsoft), and government efforts (UN’s Girls in ICT Day, European Commission’s Women in Digital Scoreboard). While women face multiple barriers to entering technology, we focus on two labor market frictions that short-term interventions can directly address. [^2] The gender gap in technology has been attributed to discrimination in hiring and promotion (Reuben et al., 2014; Moss-Racusin et al., 2012), work-family conflict and child penalties (Goldin, 2014; Kleven et al., 2019; Bertrand et al., 2010), and underrepresentation in STEM education pipelines (Bayer and Rouse, 2016; Carrell et al., 2010). We focus on network and information barriers because short-term labor market interventions can directly address these constraints, unlike structural factors requiring longer-term policy responses. First, research across multiple settings has found that women have weaker professional networks than men, limiting access to job opportunities (Greguletz et al., 2019; Forret and Dougherty, 2004). Second, gender stereotypes have been shown to lead employers to discount women’s technical abilities, creating information frictions that credible skill signals could reduce (Cheryan et al., 2013; Del Carpio and Guadalupe, 2022; Murciano-Goroff, 2022). Understanding which barriers pose the greatest impediment, and for whom, is critical for designing effective policy. If network barriers dominate, interventions should expand professional connections; if information frictions dominate, scalable credentialing programs may suffice. Alternatively, different populations may face different primary barriers, necessitating multiple program types and targeted assignment rules. We find evidence that both network deficits and information frictions matter, but for different populations: network deficits primarily constrain women outside major urban centers and those with caregiving responsibilities, while information frictions primarily constrain early-career women with limited work history.
@@
We evaluate two interventions implemented in Poland by Dare IT, a social impact organization, each targeting one of these frictions. Mentoring pairs participants with senior female technology professionals for four months of one-on-one sessions. While sessions are personalized, mentors consistently emphasize networking: facilitating introductions to hiring managers, expanding mentees’ professional connections, and teaching participants how to effectively use these connections. Challenges is a scalable online program (\$15 per participant) we designed during this research to address skill signaling barriers. The program helps participants build portfolios demonstrating technical proficiency through six bi-weekly industry-designed assignments, culminating in complete projects, such as mobile app designs or React applications. Participants later reference these projects in job applications. [^3] Program development began with structured interviews with hiring managers at twelve Polish technology firms, who uniformly cited the absence of credible practical experience signals in female applicants’ profiles. We designed Challenges to be delivered entirely online, centered on peer interaction rather than scarce mentors, and based on realistic business cases to ensure employer credibility. The program has been replicated across multiple specializations (software testing, cloud computing, UI design) and delivered in Ukrainian for refugees. Both programs were substantially oversubscribed, allowing us to randomize access and measure employment outcomes via LinkedIn profiles over twelve months. [^4] We define technology jobs as positions in technology firms (excluding finance, legal, HR, and accounting roles) or technical positions elsewhere (software development, IT support, data analytics). UX and front-end programming roles are classified as technology jobs regardless of firm type.
@@
Both programs generate large employment effects. Mentoring increases technology employment by 15 percentage points (SE 5.7, from a 46% control baseline); Challenges by 11 percentage points (SE 4.8, from 27%). Consistent with distinct mechanisms, Mentoring increases LinkedIn connections by 52.7 (SE 20.8), a 23% increase, while Challenges shows no detectable network expansion.
@@
The programs benefit different populations. Professional experience reveals the starkest contrast: Challenges generates 28 percentage points larger effects for participants with under five years of experience compared to those with longer careers, while Mentoring shows the opposite pattern with 20 percentage points larger effects for experienced workers. This indicates that early-career women benefit from portfolio signals that compensate for limited work history, while experienced women benefit from network expansion that leverages their accumulated professional standing. Geographic heterogeneity supports the network mechanism: Mentoring generates 26 percentage point effects for women in small towns compared to only 3 percentage points for Warsaw residents (difference: 23 percentage points, SE 14), while Challenges shows no significant geographic heterogeneity (difference: -2 percentage points, SE 12). This pattern is consistent with network barriers binding tightly outside major urban centers where local technology ecosystems are sparse, while online portfolio signals function equally regardless of location. These differential patterns demonstrate that both network barriers and information frictions constrain technology employment, but for distinct populations defined by career stage or geography.
@@
The differential treatment effects create an allocation problem when program capacity is limited. Both programs remain substantially oversubscribed (5:1 applicant-to-spot ratio for Mentoring; all 300 spots filled in four hours for Challenges), requiring selection among applicants. We evaluate counterfactual assignment policies using the methods of Athey and Wager (2021) and Zhou et al. (2023), estimating which applicants should receive which program to maximize aggregate outcomes. A targeted policy assigning 15% of applicants to each program based on predicted treatment effects achieves treatment effects 86% larger than randomly assigning from the pool of all applicants. This gain comes from two margins: selecting individuals with the highest treatment effects and matching individuals to their optimal program type. The matching component generates value even at scale: when capacity expands to 50% per program (universal coverage), targeted allocation maintains outcomes 23% larger than random assignment. This demonstrates that complementarity between individual characteristics and program mechanisms creates allocation gains beyond simple capacity expansion. Treatment effect heterogeneity enables improved matching even when all applicants can be served.
In Mentoring, experienced female technology professionals select mentees from a pre-screened pool based on application materials: cover letters, resumes, pre-recorded videos, and registration surveys. Mentors vary systematically in whom they select and what outcomes they achieve. Mentors with managerial experience or above-median career tenure select applicants with weaker baseline employment prospects and have stronger treatment effects: a pattern consistent with identifying "high-potential, high-barrier" candidates rather than those already likely to succeed. Less experienced mentors favor applicants with stronger baseline employability, selecting candidates who would likely find technology jobs regardless of program participation. If effective selection heuristics can be learned through experience, data-driven tools may offer a path to improving targeting among less experienced selectors—a possibility we examine next.
@@
We find that they can. Algorithmic allocation using only six observable characteristics (education, location, age, professional experience, degree level, domain expertise) achieves treatment effects 63% larger than random assignment and 11% than average mentor selections. The two approaches achieve substantial concordance (70%): the algorithm largely validates expert choices while identifying additional high-benefit candidates. Despite mentors’ rich soft information, systematic patterns in treatment response remain that simple demographic targeting can exploit. This result extends recent work showing algorithms can complement expert decisions in domains from bail (Kleinberg et al., 2018) to medical diagnosis (Mullainathan and Obermeyer, 2022), demonstrating that even sophisticated expert judgment leaves room for data-driven improvement when treatment effects vary systematically with observable characteristics.
Our focus on small, custom-designed programs limits external validity but enables deeper investigation of mechanisms. Mentoring programs are almost always small due to capacity constraints. [^5] Our 150 participants is comparable to Dennehy and Dasgupta (2017) and exceeds the 30–40 per cohort in Blau et al. (2010). Working with each mentor, we implemented a paired randomization design that allows us to measure both selection and treatment effects—observing not only program impacts but how mentor choices vary with mentor experience. We developed Challenges during this research and evaluate its first cohort in this paper; the program has since scaled to multiple technology domains (software testing, cloud computing, UI design) and languages of instruction, suggesting the model extends beyond our pilot context.
@@ 2 Literature Review
This paper contributes to the literature on treatment effect heterogeneity and optimal policy assignment. Building on recent advances in policy learning using machine learning methods (Athey and Imbens, 2017; Kitagawa and Tetenov, 2018; Athey and Wager, 2021), we show that data-driven targeting can substantially improve program allocation decisions in active labor market policy. Theoretical work demonstrates that when treatment effects vary across individuals and program capacity is limited, policies that assign treatment based on predicted treatment effects can substantially improve welfare relative to random allocation or allocation based on baseline characteristics (Kitagawa and Tetenov, 2018; Bhattacharya and Dupas, 2012). Empirical applications of these methods have demonstrated value across diverse domains: customer retention (Ascarza, 2018), marketing (Hitsch et al., 2024), development programs (Haushofer et al., 2025), agriculture (Kasy and Sautmann, 2021), and medical treatment (Inoue et al., 2023).
@@
We further contribute to the growing literature examining whether algorithmic decision-making can complement or improve upon expert human judgment. Recent work demonstrates that machine learning algorithms can match or exceed expert performance in domains from bail decisions (Kleinberg et al., 2018) to medical diagnosis (Mullainathan and Obermeyer, 2022), often by identifying systematic patterns that experts miss or cannot consistently apply. Similar to findings in online labor markets where revealing information about worker quality improves matching outcomes (Pallais, 2014), we show that observable characteristics predict treatment effects systematically.
@@
We demonstrate that data-driven targeting can improve upon expert human judgment. Mentor-based selection achieves treatment effects 48% larger than random allocation. Yet algorithmic targeting achieves an additional 11% improvement. The 70% concordance between methods indicates complementarity: algorithms largely validate expert choices while identifying additional high-benefit candidates, demonstrating that systematic patterns in treatment response exist even after expert selection captures most available heterogeneity.
@@
This paper contributes to the literature on mentoring effectiveness in improving career outcomes. Several influential papers focus on mentoring in academic contexts and show that mentorship increases sense of belonging among female engineering students (Dennehy and Dasgupta, 2017), youth’ learning outcomes (Resnjanskij et al., 2024), and tenure rates (Ginther et al., 2020), publications, and federal grants for women in research careers (Gardiner et al., 2007; Blau et al., 2010). [^6] The impact of mentoring on career outcomes has also been studied using observational data. Ragins and Cotton (1999) survey over 600 individuals and find that mentored respondents have better career outcomes than non-mentored respondents, with effects varying by the gender composition of the mentoring pair. Kammeyer-Mueller and Judge (2008) conduct a meta-analysis concluding that mentoring improves job and career satisfaction. Mentoring programs may operate partly through expanding professional networks, which play a critical role in labor market outcomes (Granovetter, 1973; Ioannides and Loury, 2004; Pallais and Sands, 2016). Athey et al. (2000) develop a formal theoretical model of the dynamics of mentoring and diversity. We provide randomized trial evidence on mentoring effectiveness in labor market transitions outside academic settings, demonstrating large effects on technology employment for women career-changers. Crucially, we provide experimental evidence that mentoring causally expands professional networks: Mentoring increases participants’ LinkedIn connections by 52.7 (SE 20.8), a 23% increase from baseline, while the non-network-based Challenges program shows no statistically significant effect. We further examine the evaluative selection mechanism inherent in mentor-based programs, where mentors exercise discretion in choosing mentees, and demonstrate that this selection generates substantial treatment effect heterogeneity with more experienced mentors systematically selecting mentees with higher treatment effects.
To our knowledge, we are the first to evaluate mentoring effectiveness using a paired randomization design within mentor-selected candidates. This approach offers three advantages over standard randomization from an applicant pool. First, it allows us to estimate causal treatment effects while preserving the program’s natural selection structure—mentors choose whom they want to work with, avoiding the forced assignments that standard randomization would impose and that mentors might resist. Second, by observing mentor choices, we can study how selection varies with mentor characteristics and whether experienced mentors develop effective selection heuristics. Third, because we observe outcomes for non-selected applicants alongside selected participants randomized to control, we can directly compare these groups, which is not possible in prior mentoring RCTs, which randomize only within pre-screened applicant pools.
@@
We contribute to the extensive literature evaluating active labor market training programs. Card et al. (2018) conduct a comprehensive meta-analysis of over 200 studies, finding that traditional training programs aimed at unemployed workers exhibit near-zero short-term employment effects, with modest positive impacts emerging only 2-3 years post-completion. [^7] See also Heckman et al. (1999), Card et al. (2010), and Crépon and van den Berg (2016) for reviews of ALMP effectiveness. Programs incorporating on-the-job skill development tend to outperform classroom-based alternatives (Sianesi, 2008; Lechner and Gerfin, 2000), though even intensive apprenticeships show mixed results (Biewen et al., 2014). Fein and Hamadyk (2018) show that YearUp, a sectoral training program targeting disadvantaged youth, generates 40% earnings gains. However, recent evidence suggests that programs targeting skill signaling and information frictions generate substantially larger employment gains than traditional classroom training (Abebe et al., 2021; Bassi and Nansamba, 2022; Carranza et al., 2022; Athey and Palikot, 2024). Building on the theoretical work of Spence (1978), this literature shows that credentials demonstrating capability can significantly boost earnings (Tyler et al., 2000; Hadavand et al., 2018) and generate market-level benefits when employers face uncertainty about worker quality (Pallais, 2014). Recent evidence further suggests that training effectiveness may depend on complementary interventions: Batista et al. (2022) find that business training for female microentrepreneurs generates significant profit increases only when combined with mobile money access, highlighting that releasing multiple constraints simultaneously may be necessary for training programs to translate into economic gains.
@@
We provide experimental evidence on training program effectiveness in technology sector transitions, a context involving higher-skilled, career-changing women rather than the unemployed workers typically studied. We document substantial short-term effects (visible within twelve months) for both programs, aligning with the signaling literature rather than the null findings in Card et al. (2018)’ meta-analysis. We directly compare two program modalities: personalized one-on-one mentoring versus scalable online project-based training. This head-to-head comparison reveals complementarity rather than substitutability at a policy-portfolio level: the programs serve different populations and generate heterogeneous treatment effects that enable targeted assignment. Our subjects are women seeking technology roles who can utilize skill signaling to counter gender-based statistical discrimination.
@@
Finally, this paper adds to the literature on the gender gap in the technology sector. The persistent gender wage gap reflects both occupational segregation and within-occupation differences (Juhn and McCue, 2017). Various reasons for low women’s participation in the technology sector have been studied, including women’s tendency to under-report programming skills (Murciano-Goroff, 2022) and the impact of occupational stereotypes (Cheryan et al., 2013; Del Carpio and Guadalupe, 2022). Strategies proposed to counter these barriers include increasing role model visibility (Correll and Mackenzie, 2016) and shifting perceptions of technology roles. We contribute to this literature by demonstrating how mentoring and credentialing programs can enhance women’s success rates in obtaining technology jobs.